OpenStudy (anonymous):

Milo bought 4 1/2 pounds of red apples and 5 5/6 green apples. How many lbs of apples did he buy in all? Write an expression that models the problem. Help me please

OpenStudy (anonymous):

how does this sound? 4 + 5 = 9 3/6 + 5/6 = 8/6 = 1 2/6 /1 1/3 9 + 1 1/3 = 10 1/3 lbs of apples

OpenStudy (anonymous):

did already u write the Expression

OpenStudy (anonymous):

I wrote it just like that - What do you think? Is it right?

OpenStudy (anonymous):

I am on another problem can you help me? combine like terms 11 + 8x - 11 x + 13

OpenStudy (anonymous):

would it be 24 - 3x ?

OpenStudy (texaschic101):

11 + 8x - 11x + 13 --- combine like terms -3x + 24...or 24 - 3x....you are correct